In Compton scattering, the incoming and scattered photons have a different wavelength. If the outgoing photon has a wavelength which is 12% more than the incoming, what is the ratio of the outgoing photon energy to the incoming photon energy (E subscript text out end text end subscript over E subscript text in end text end subscript)? (2 s.f.)
1
Expert's answer
2020-05-29T09:51:50-0400

Given

Wavelength of outgoing photon is "\lambda _{out}=(0.12+1)\lambda _{in}=1.12 \lambda _{in}"

The ratio of the outgoing photon energy to the incoming photon energy is

"\frac{E_{out}}{E_{in}}=\frac{\frac{hc}{\lambda _{out}}}{\frac{hc}{\lambda _{in}}}"

"\frac{E_{out}}{E_{in}}=\frac{\lambda _{in}}{\lambda _{out}}"

"=\frac{\lambda _{in}}{1.12\lambda _{in}}"

"=\frac{1}{1.12}=0.89"

Hence, the ratio of the outgoing photon energy to the incoming photon energy is"\frac{E_{out}}{E_{in}}=\frac{1}{1.12}=0.89" .